Requirement 1h. Compute the earnings per share of common stock for 2018 and 2017. Begin by selecting the formula to compute the earnings per share of common stock.

Earnings per share

=

(Net income - Preferred dividends) / Weighted average number of common shares outstanding

​Now, compute the earnings per share for 2018 and 2017.

​(Round your answers to the nearest​ cent, $X.XX.)

2018:

2017:

Banfield, Inc.

Comparative Balance Sheet

December 31, 2018 and 2017

2018

2017

2016*

Assets

Current Assets:

Cash

$90,000

$89,000

Accounts Receivables, Net

111,000

114,000

$105,000

Merchandise Inventory

149,000

162,000

210,000

Prepaid Expenses

16,000

8,000

Total Current Assets

366,000

373,000

Property, Plant, and Equipment, Net

220,000

173,000

Total Assets

$586,000

$546,000

$600,000

Liabilities

Total Current Liabilities

$227,000

$246,000

Long-term Liabilities

120,000

97,000

Total Liabilities

347,000

343,000

Stockholders' Equity

Preferred Stock, 5%

110,000

110,000

Common Stockholders' Equity, no par

129,000

93,000

88,000

Total Liabilities and Stockholders' Equity

$586,000

$546,000

* Selected 2016 amounts

Banfield, Inc.

Comparative Income Statement

Years Ended December 31, 2018 and 2017

2018

2017

Net Sales Revenue

$464,000

$428,000

Cost of Goods Sold

241,000

218,000

Gross Profit

223,000

210,000

Operating Expenses

138,000

136,000

Income From Operations

85,000

74,000

Interest Expense

14,000

15,000

Income Before Income Tax

71,000

59,000

Income Tax Expense

20,000

26,000

Net Income

$51,000

$33,000
Answer #1

1. Solve the earing per share as formula given below :

formula :

  earning per share = ( net income - prefered dividend ) / out standing share

A. earning per share = ( 500000 - 180000) 250000= 1.28 per share

B . earning per share = ( 850000- 90000) 200000) =3.80 per share